Jose has $30,000 that he puts into two accounts. The first account pays 8% simple interest and the second account pays 5% simple interest. At the end of one year, he has earned a total of $2250 in interest. How much did he have in each account?
Answer by jorel1380(3719)

.08n+.05(30000-n)=2250
8n+5(30000-n)=225000
8n+150000-5n=225000
3n=75000
n=25000
30000-n=5000
$25000 was invested at 8%, and $5000 was invested at 5%..